# Quantum Advantage in Private Multiple Hypothesis Testing

### v1

@@ I Introduction
In the past decade, the concept of LDP has been extended to quantum systems, which we call quantum LDP (QLDP) [21, 22, 23, 24]. The most relevant works with our work were done by Yoshida and Hayashi [22, 25]: Yoshida and Hayashi [22] showed that there is no quantum advantage in private statistical inference if original classical data is binary and the utility measure satisfies the DPI. This impossibility result was proved by showing that every $\epsilon$-QLDP mechanism is essentially classical, i.e., it is a composition of an $\epsilon$-LDP mechanism and a classical-quantum (CQ) channel, if classical data is binary ($\epsilon$ is a positive real number representing the limit of acceptable privacy leakage). Recently, Yoshida [25] showed that there is a quantum advantage in a certain optimization problem under privacy constraint.
In this paper, we demonstrate a quantum advantage in a certain operational scenario, a private multiple hypothesis testing where the original data is classical. The privacy constraint is set to QLDP, and the utility is chosen as the pairwise-minimum Chernoff information because it characterizes the rate of the optimal average error probability of testing as proved by Li [26]. Concerning some class of hypotheses that we call smoothed point masses, we demonstrate that the quantum advantage with respect to the optimal PUT exists. This result is proved by deriving an upper bound of the optimal classical PUT that is tight for some cases, and proposing a certain QLDP mechanism which achieves a better PUT than such an upper bound. To derive the upper bound of the optimal classical PUT, we exploit extremal mechanisms proposed by Kairouz et al. [13], and block design mechanisms proposed by Park et al. [27] shows the tightness of the bound for some cases. The proposed QLDP mechanism is a composition of a CQ channel whose outputs are pure states, and a depolarizing channel. Here, pure states and the level of depolarizing noise are suitably chosen to maximize the utility at some high privacy regime and for some hypotheses, while satisfying the QLDP constraint. The chosen pure states are states corresponding to a symmetric informationally complete positive operator-valued measure (SIC-POVM) [28, 29].
@@ II Problem formulation
For $v=2$, there is no quantum advantage, i.e., $M_{\mathrm{Q}}(2,\epsilon;\mathbf{H})=M_{\mathrm{C}}(2,\epsilon;\mathbf{H})$ for all hypotheses $\mathbf{H}$ and $\epsilon>0$. This fact is a direct consequence of the results by Yoshida and Hayashi [22]: every $\epsilon$-QLDP mechanism is a sequential composition of an $\epsilon$-LDP mechanism and a CQ channel, or simply, an essentially classical $\epsilon$-QLDP mechanism. Because the Chernoff information satisfies the DPI [31, Chap. 4.4.1], there is no quantum advantage if $v=2$.
@@ III-A Optimal classical PUT
The proof of the proposition consists of the converse part and the achievability part. Concerning the converse part, we show that it is enough to consider extremal mechanisms [13], and upper-bound the utility of an extremal mechanism using some techniques used in the proof of [25, Lem. 5.3]. For the achievability part, the utility of a block design mechanism [27] is shown to meet the converse bound at $\eta=1$. The proof of (19) is based on Taylor’s approximation.
@@ III-A1 Converse
Note that for probability vectors $(p,p^{\prime})$, $\sum\limits_{y}p_{y}^{s}(p^{\prime})_{y}^{1-s}$ is jointly concave in $(p,p^{\prime})$ for $s\in[0,1]$ [31, Prop. 4.8]. This implies that Chernoff information satisfies the DPI, and so does $M^{\eta}$. Therefore, Lemma 4 tells that $M_{\mathrm{C}}^{\eta}(v,\epsilon)=M(q)$ for some extremal $\epsilon$-LDP mechanism $q=S^{(v)}\mathrm{diag}(\theta)$. Thus,
@@
Now, we use similar techniques to those in the proof of [25, Lem. 5.3]. Based on the above argument, we get
@@ III-A2 Achievability
An $r$-regular, $k$-uniform, $\lambda$-pairwise balanced BIBD is conventionally written as $(v,k,\lambda)$-design, because $r,b$ can be determined by $(v,k,\lambda)$. Whenever $k\notin\{0,1,v\}$, the following identities should hold [32, Thm. 1.8, 1.9]:
@@ III-B Proposed QLDP mechanism
In this subsection, we consider a class of QLDP mechanisms which are sequential compositions of a CQ channel whose outputs are pure states, and a depolarizing channel. Precisely, we consider a privacy mechanism $Q$ in the following form:
@@
where each of |$\ket{\psi_{x}}$ is a unit vector in $\mathcal{H}_{d}$ and $\mu\in[0,\frac{d}{d-1}]$. [^1] The depolarizing channel with parameter $\mu$ is a valid quantum channel (trace-preserving and completely positive) if and only if $\mu\in[0,\frac{d^{2}}{d^{2}-1}]$. Though, we allow $\frac{d^{2}}{d^{2}-1}<\mu\leq\frac{d}{d-1}$ since $Q$ a valid classical-quantum channel if and only if $\mu\in[0,\frac{d}{d-1}]$. It can be easily observed that $Q_{x}$ is positive semidefinite and $\mathrm{Tr}(Q_{x})=1$. This class of privacy mechanisms with $\mu\in(0,1]$ was previously considered by Yoshida [25]. We further choose suitable depolarizing parameter $\mu$ and pure states {|$\{\ket{\psi_{x}}\}_{x\in[v]}$ which maximize the utility at $\eta=1$ and $\epsilon\approx 0$, while ensuring the $\epsilon$-QLDP constraint. The chosen privacy mechanism shows a quantum advantage at $13\leq v\leq 9,\eta=$ 1, and $\epsilon\approx 0$. If $v=4$ or $9$, then the advantage can be extended to all $\eta$ close to $1$ by the continuity of $M^{\eta}$ and $\bar{M}_{\mathrm{C}}^{\eta}$.
We first derive a sufficient and necessary condition for such a privacy mechanism to be an $\epsilon$-QLDP mechanism, whose proof is similar to that of [25, Lem. 5.5].
@@
According to Proposition 7, we choose a set of pure states as a subset of a SIC-POVM in Definition 3. Also, it can be seen that smaller $d$ gives a better utility. Concerning this point, the number of pure states {|$\{\ket{\psi_{x}}\}$ in $\mathcal{H}_{d}$ with constant $c_{x,x^{\prime}}$ is at most $d^{2}$ [29, Thm. 2.28], and SIC-POVM in Definition 3 is such a maximal set. The existence of SIC-POVM is known up to $d\leq 151$, and it has been conjectured that SIC-POVM exists for all $d$ [34, 29, 28]. Combining with these facts and Propositions 3 and 7, we can show quantum advantage at $\eta=1$ as in the following corollary whose proof is straightforward.

### v2

@@ I Introduction
In the past decade, the concept of (L)DP has been extended to quantum systems, which we call quantum (L)DP (Q(L)DP) [21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35]. Among them, [26, 29, 33, 34] considered private hypothesis testing under Q(L)DP constraint, and provided bounds on the optimal PUT with respect to some utility measures related to error probabilities. However, these works did not show whether a quantum advantage exists or not. Regarding quantum advantage, which is the focus of this work, Yoshida and Hayashi showed the following results [23, 35]: [23] showed that there is no quantum advantage in private statistical inference if original classical data is binary and the utility measure satisfies the DPI. This impossibility result was proved by showing that every $\epsilon$-QLDP mechanism is essentially classical, i.e., it is a composition of an $\epsilon$-LDP mechanism and a classical-quantum (CQ) channel, if classical data is binary ($\epsilon$ is a positive real number representing the limit of acceptable privacy leakage). Recently, Yoshida [35] showed that there is a quantum advantage in a certain optimization problem under privacy constraint.
In this paper, we demonstrate a quantum advantage in a certain operational scenario, a private multiple hypothesis testing where the original data is classical. The privacy constraint is set to QLDP, and the utility is chosen as the pairwise-minimum Chernoff information because it characterizes the rate of the optimal average error probability of testing as proved by Li [36]. Concerning some class of hypotheses that we call smoothed point masses, we demonstrate that the quantum advantage with respect to the optimal PUT exists. This result is proved by deriving an upper bound of the optimal classical PUT that is tight for some cases, and proposing a certain QLDP mechanism which achieves a better PUT than such an upper bound. To derive the upper bound of the optimal classical PUT, we exploit extremal mechanisms proposed by Kairouz et al. [13], and block design mechanisms proposed by Park et al. [37] show the tightness of the bound for some cases. The proposed QLDP mechanism is a composition of a CQ channel whose outputs are pure states, and a depolarizing channel. Here, pure states and the level of depolarizing noise are suitably chosen to maximize the utility at some high privacy regime and for some hypotheses, while satisfying the QLDP constraint. The chosen pure states are states corresponding to a symmetric informationally complete positive operator-valued measure (SIC-POVM) [38, 39].
@@ II Problem formulation
For $v=2$, there is no quantum advantage, i.e., $M_{\mathrm{Q}}(2,\epsilon;\mathbf{H})=M_{\mathrm{C}}(2,\epsilon;\mathbf{H})$ for all hypotheses $\mathbf{H}$ and $\epsilon>0$. This fact is a direct consequence of the results by Yoshida and Hayashi [23]: every $\epsilon$-QLDP mechanism is a sequential composition of an $\epsilon$-LDP mechanism and a CQ channel, or simply, an essentially classical $\epsilon$-QLDP mechanism. Because the Chernoff information satisfies the DPI [41, Chap. 4.4.1], there is no quantum advantage if $v=2$.
@@ III-A Optimal classical PUT
The proof of the proposition consists of the converse part and the achievability part. Concerning the converse part, we show that it is enough to consider extremal mechanisms [13], and upper-bound the utility of an extremal mechanism using some techniques used in the proof of [35, Lem. 5.3]. For the achievability part, the utility of a block design mechanism [37] is shown to meet the converse bound at $\eta=1$. The proof of (19) is based on Taylor’s approximation.
@@ III-A1 Converse
Note that for probability vectors $(p,p^{\prime})$, $\sum\limits_{y}p_{y}^{s}(p^{\prime})_{y}^{1-s}$ is jointly concave in $(p,p^{\prime})$ for $s\in[0,1]$ [41, Prop. 4.8]. This implies that Chernoff information satisfies the DPI, and so does $M^{\eta}$. Therefore, Lemma 4 tells that $M_{\mathrm{C}}^{\eta}(v,\epsilon)=M(q)$ for some extremal $\epsilon$-LDP mechanism $q=S^{(v)}\mathrm{diag}(\theta)$. Thus,
@@
Now, we use similar techniques to those in the proof of [35, Lem. 5.3]. Based on the above argument, we get
@@ III-A2 Achievability
An $r$-regular, $k$-uniform, $\lambda$-pairwise balanced BIBD is conventionally written as $(v,k,\lambda)$-design, because $r,b$ can be determined by $(v,k,\lambda)$. Whenever $k\notin\{0,1,v\}$, the following identities should hold [42, Thm. 1.8, 1.9]:
@@ III-B Proposed QLDP mechanism
In this subsection, we propose a QLDP mechanism which shows a quantum advantage. We consider a class of QLDP mechanisms which are sequential compositions of a CQ channel whose outputs are pure states, and a depolarizing channel. Precisely, we consider a privacy mechanism $Q$ in the following form:
@@
where each of |$\ket{\psi_{x}}$ is a unit vector in $\mathcal{H}_{d}$ and $\mu\in[0,\frac{d}{d-1}]$. [^1] The depolarizing channel with parameter $\mu$ is a valid quantum channel (trace-preserving and completely positive) if and only if $\mu\in[0,\frac{d^{2}}{d^{2}-1}]$. Though, we allow $\frac{d^{2}}{d^{2}-1}<\mu\leq\frac{d}{d-1}$ since $Q$ a valid classical-quantum channel if and only if $\mu\in[0,\frac{d}{d-1}]$. It can be easily observed that $Q_{x}$ is positive semidefinite and $\mathrm{Tr}(Q_{x})=1$. This class of privacy mechanisms with $\mu\in(0,1]$ was previously considered by Yoshida [35]. Compared to previous work considering a depolarizing channel as a Q(L)DP mechanism [21, 27, 29, 30, 31, 32, 34, 35], we further choose suitable depolarizing parameter $\mu$ and pure states {|$\{\ket{\psi_{x}}\}_{x\in[v]}$ which maximize the utility at $\eta=1$ and $\epsilon\approx 0$, while ensuring the $\epsilon$-QLDP constraint. The chosen privacy mechanism shows a quantum advantage at $13\leq v\leq 9,\eta=$ 1, and $\epsilon\approx 0$. If $v=4$ or $9$, then the advantage can be extended to all $\eta$ close to $1$ by the continuity of $M^{\eta}$ and $\bar{M}_{\mathrm{C}}^{\eta}$.
We first derive a sufficient and necessary condition for such a privacy mechanism to be an $\epsilon$-QLDP mechanism, whose proof is similar to that of [35, Lem. 5.5].
@@
According to Proposition 7, we choose a set of pure states as a subset of a SIC-POVM in Definition 3. Also, it can be seen that smaller $d$ gives a better utility. Concerning this point, the number of pure states {|$\{\ket{\psi_{x}}\}$ in $\mathcal{H}_{d}$ with constant $c_{x,x^{\prime}}$ is at most $d^{2}$ [39, Thm. 2.28], and SIC-POVM in Definition 3 is such a maximal set. The existence of SIC-POVM is known up to $d\leq 151$, and it has been conjectured that SIC-POVM exists for all $d$ [44, 39, 38]. Combining with these facts and Propositions 3 and 7, we can show quantum advantage at $\eta=1$ as in the following corollary whose proof is straightforward.